OpenStudy (anonymous):

2x^3 y^4 z^3 a^2 (over (/)) 4x^2 y^5 za Equals?

OpenStudy (anonymous):

I'm really not sure what I'm supposed to do with this problem. I attached the equation (I couldn't put it in the question title).

OpenStudy (anonymous):

\(\Large{\dfrac{2x^3y^4z^3a^2}{4x^2y^5za}}\) Start by canceling.

OpenStudy (anonymous):

When dividing, subtract the exponents. \(x^3 \div x^2 = \x^{3-2} = x\) \(\Large{\dfrac{2\cancel{x^3}(x)y^4z^3a^2}{4\cancel{x^2}y^5za}}\)